OpenStudy (anonymous):

a stream is a mile wide and flowing 2 mph. Kristi believes she is rowing directly across the stream in her dingy at 6 mph. How fast will Kristi actually travel? How far down stream will she land? I just need to know how to set it up.

OpenStudy (anonymous):

So I would draw a line heading north representing the vector 6mph. Then draw a line at the end of the first line towards east representing the vector 2mph. Then connect the two ends of the lines to form an inverted right triangle. You need to get the angle of the point closest to you and you can get that by\[\tan inverse (2/6) = \angle\]

OpenStudy (anonymous):

to get the speed at which she is actually traveling, use pythagorean theorem to get the resultant vector \[\sqrt(a^2+b^2)\]

OpenStudy (anonymous):

to get the distance she would have travelled downstream, use \[\sin \theta = y/resultant vector\]

OpenStudy (anonymous):

oops check that \[\tan \theta = y/1 mile\]